# CAST: Achieving Stable LLM-based Text Analysis for Data Analytics

### v1

@@ Abstract
Text analysis of tabular data relies on two core operations: summarization for corpus-level theme extraction and tagging for row-level labeling. A critical limitation of employing large language models (LLMs) for these tasks is their inability to meet the high standards of output stability demanded by data analytics. To address this challenge, we introduce CAST (Consistency via Algorithmic Prompting and Stable Thinking), a framework that enhances output stability by constraining the model’s latent reasoning path. CAST combines (i) Algorithmic Prompting to impose a procedural scaffold over valid reasoning transitions and (ii) Thinking-before-Speaking to enforce explicit intermediate commitments before final generation. To measure progress, we introduce CAST-S and CAST-T, stability metrics for bulleted summarization and tagging, and validate their alignment with human judgments. Experiments across publicly available benchmarks on multiple LLM backbones show that CAST consistently achieves the best stability among all baselines, improving Stability Score by up to 16.2%, while maintaining or improving output quality.
@@
Jinxiang Xie1thanks: The work was done during the author’s internship at Microsoft., Zihao Li211footnotemark: 1, Wei He311footnotemark: 1, Rui Ding4thanks: Corresponding author: juding@microsoft.com., Shi Han4, Dongmei Zhang4, 1Nanjing University, 2Tsinghua University, 3Peking University, 4Microsoft Research, Correspondence: xiejinxiang@smail.nju.edu.cn, juding@microsoft.com
@@ 1 Introduction
These operators are not only expressive but also composable. For example, “summarizing negative reviews” can be implemented by (i) tagging sentiment, (ii) filtering the negative subset, and (iii) summarizing the filtered corpus. This composability suggests a small set of principled text operators can support a wide range of analytical queries.
Large language models (LLMs) are natural candidates to operationalize TADA, since a single model can execute diverse text analyses via a natural-language query. In this paradigm, many classical tasks (sentiment analysis, keyword extraction) become instances of tagging, parameterized by the query. More broadly, TADA shifts system design from a method-centric multi-model pipeline to a query-centric LLM operator:
@@
The stability requirement. Despite this promise, TADA exposes a critical misalignment between the probabilistic nature of LLM generation and the deterministic requirements of data analytics. In creative settings, diversity is desirable. In TADA, however, stability is a functional necessity: once tags or themes are materialized as columns, they become keys for grouping and aggregation. If the same review is labeled as “Customer Service” in one run and “Support Team” in another, downstream results can change, undermining reproducibility and trust Atil et al. (2024); Croxford et al. (2025).
@@ Acknowledgements
We are grateful to Yue Wang and Liu Ye for their insightful feedback. We thank the anonymous reviewers for their constructive comments. Rui Ding is the corresponding author.
@@ A.3 Large Language Model Integration
Our implementation integrates multiple LLM providers through a unified API interface, supporting models including GPT-5, GPT-5.2 from OpenAI, DeepSeek-V3.2 and DeepSeek-V3.2-Exp (Noted as D.S.-V3.2-Exp) and Qwen3 series. Each provider is configured with appropriate timeout settings (300 seconds) and error handling mechanisms to ensure robust execution across different API constraints and rate limits. Model parameters are set consistently: temperature is 0, seed is 42.
@@ C.1 Details of CAST-S for Summarization Stability
Notes on judge robustness.
LLM-as-a-judge can exhibit known biases such as positional and verbosity bias Ye et al. (2024). CAST-S mitigates these issues by (i) relying on pairwise matching rather than absolute grading, (ii) using two independent judges and averaging, and (iii) explicitly incorporating an ordering component that is computed algorithmically.
@@ D.1 Additional Results from Different LLMs
These supplementary results confirm the consistent performance of CAST across different model scales.
@@
[…]

### v2

@@ Abstract
Text analysis of tabular data relies on two core operations: summarization for corpus-level theme extraction and tagging for row-level labeling. A critical limitation of employing large language models (LLMs) for these tasks is their inability to meet the high standards of output stability demanded by data analytics. To address this challenge, we introduce CAST (Consistency via Algorithmic Prompting and Stable Thinking), a framework that enhances output stability by constraining the model’s latent reasoning path. CAST combines (i) Algorithmic Prompting to impose a procedural scaffold over valid reasoning transitions and (ii) Thinking-before-Speaking to enforce explicit intermediate commitments before final generation. To measure progress, we introduce CAST-S and CAST-T, stability metrics for bulleted summarization and tagging, and validate their alignment with human judgments. Experiments across publicly available benchmarks on multiple LLM backbones show that CAST consistently achieves the best stability among all baselines, improving Stability Score by up to 16.2%, while maintaining or improving output quality. [^1] Code, prompts, and data are available at https://github.com/jxtse/CAST-text-analysis.
@@
Jinxiang Xie1thanks: These authors contributed equally. [^3] The work was done during the authors’ internship at Microsoft., Zihao Li211footnotemark: 1333The work was done during the authors’ internship at Microsoft., Wei He3333The work was done during the authors’ internship at Microsoft., Rui Ding4thanks: Corresponding author., Shi Han4, Dongmei Zhang4, 1Nanjing University, 2Tsinghua University, 3Peking University, 4Microsoft Research, Correspondence: xiejinxiang@smail.nju.edu.cn, juding@microsoft.com
@@ 1 Introduction
These operators are not only expressive but also composable. For example, “summarizing negative reviews” can be implemented by (i) tagging sentiment, (ii) filtering the negative subset, and (iii) summarizing the filtered corpus. This composability suggests a small set of principled text operators can support a wide range of analytical queries. Importantly, tagging and summarization are not independent tasks, but instead form mutually reinforcing stages in a unified analytical pipeline. Tagging produced at the row level provides the structured keys upon which summarization can condition, while the themes identified during summarization can, in turn, serve as a controlled vocabulary for subsequent tagging passes. By ensuring stability in both operators, CAST prevents error propagation across this “Tag–Filter–Summarize” pipeline, which is the dominant workflow in production TADA systems.
LLMs are natural candidates to operationalize TADA, since a single model can execute diverse text analyses via a natural-language query. In this paradigm, many classical tasks (sentiment analysis, keyword extraction) become instances of tagging, parameterized by the query. More broadly, TADA shifts system design from a method-centric multi-model pipeline to a query-centric LLM operator:
@@
The stability requirement. Despite this promise, TADA exposes a critical misalignment between the probabilistic nature of LLM generation and the deterministic requirements of data analytics. In creative settings, diversity is desirable. In TADA, however, stability is a functional necessity: once tags or themes are materialized as columns, they become keys for grouping and aggregation. If the same review is labeled as “Customer Service” in one run but “Support Team” in another, downstream results can change, undermining reproducibility and trust Atil et al. (2024); Croxford et al. (2025).
@@ 2 Related Work
Stability and Reliability of LLMs. A growing body of work examines the stability and reliability of LLM outputs from complementary perspectives. Mahaut et al. (2024) show that confidence estimates for factual claims are often unstable across semantically equivalent inputs, revealing fragility in the model’s parametric knowledge. Cheng et al. (2025) propose a unified robustness framework based on distance mapping distortion to evaluate sample-level stability under input perturbations without modifying model parameters. Zhao et al. (2024) provide a comprehensive survey on the reliability of LLMs under misinformed and unconventional prompts, covering distributional perturbation analysis, sentiment analysis, and categorization. These works primarily address input-side stability (robustness to perturbation) or factual confidence. In contrast, our work targets output-side stability: ensuring that identical inputs and decoding settings yield consistent structured outputs across independent runs, a requirement specific to the deterministic workflows of data analytics.
@@ Acknowledgements
We are grateful to Yue Wang, Liu Ye and Yifan Chen for their insightful feedback. We thank the anonymous reviewers for their constructive comments.
@@
Regarding efficiency, CAST’s structured prompt and mandatory intermediate-state generation introduce a modest token overhead compared to vanilla prompting. As shown in the output JSON schema (Appendix A.1), the intermediate states (domain, topic schema, clusters) are condensed into concise key-value pairs, empirically adding only a few tens of tokens per call. In our experiments (Table 2, Panel B), CAST’s latency increase over baselines is marginal (e.g., 6.60s vs. 6.06s on Gemini-3-Flash), especially when contrasted with Self-Consistency which requires $3\times$ or more compute. For large-scale deployments, CAST integrates naturally with batch-processing and MapReduce architectures: the framework operates on micro-batches (e.g., 200 rows) in the Map phase and aggregates partial results in a Reduce pass, thereby avoiding context-window saturation.
@@ A.3 Large Language Model Integration
Our implementation integrates multiple LLM providers through a unified API interface, supporting models including GPT-5, GPT-5.2 from OpenAI, DeepSeek-V3.2 and DeepSeek-V3.2-Exp (Noted as D.S.-V3.2-Exp), Qwen3 series, Claude-Sonnet-4.6 from Anthropic, and LLaMA-4-Maverick from Meta. Each provider is configured with appropriate timeout settings (300 seconds) and error handling mechanisms to ensure robust execution across different API constraints and rate limits. Model parameters are set consistently: temperature is 0, seed is 42.
@@ C.1 Details of CAST-S for Summarization Stability
Notes on ordering rules and judge robustness.
A natural concern is whether ordering constraints hard-coded in the CAST prompt (e.g., descending topic weight, alphabetical tie-breaking) create a circularity that inflates CAST-S scores. We argue that these rules reflect a functional requirement of the TADA domain rather than a metric artifact: in tabular analytics, when a dashboard refreshes and summary bullets appear in a different order, users perceive this as system instability. Therefore, measuring and imposing order consistency via Kendall’s $\tau$ is integral to the task definition. Furthermore, CAST-S does not simply reward formatting compliance; its semantic component ($S_{sem}$) accounts for the majority of the score ($\alpha=0.9$), ensuring that content agreement dominates. Regarding LLM-as-a-judge biases such as positional and verbosity bias Ye et al. (2024), CAST-S mitigates these by (i) relying on pairwise matching rather than absolute grading, (ii) using two independent judges and averaging, and (iii) incorporating an ordering component computed algorithmically rather than by the judge.
@@ D.1 Additional Results from Different LLMs
These supplementary results confirm the consistent performance of CAST across different model scales and architectures, including models from Anthropic (Claude-Sonnet-4.6) and Meta (LLaMA-4-Maverick), demonstrating that CAST provides a universal stabilizing layer across all major LLM families.
@@
GPT-5 Zero-shot CoT $8.33\pm 0.56$ $8.55\pm 0.77$ $20.44\pm 3.74$ $67.82\pm 21.72$
Few-shot CoT $8.86\pm 0.70$ $8.62\pm 0.79$ $9.64\pm 1.78$ $62.13\pm 23.68$
Self-Consistency $9.01\pm 0.32$ $8.66\pm 0.58$ $52.02\pm 18.85$ $166.26\pm 66.13$
AP-Only $9.17\pm 0.72$ $8.87\pm 0.91$ $33.39\pm 12.15$ $83.22\pm 23.55$
TbS-Only $8.86\pm 0.78$ $8.73\pm 0.85$ $28.15\pm 7.01$ $100.81\pm 29.15$
CAST (AP+TbS) 9.27 $\pm$ 0.82 8.91 $\pm$ 0.93 $33.28\pm 5.81$ $93.50\pm 34.76$
Qwen3-Next Zero-shot CoT $8.60\pm 0.70$ $8.38\pm 0.79$ $11.08\pm 3.07$ $187.23\pm 57.66$
Few-shot CoT $8.55\pm 0.84$ $8.44\pm 0.39$ $22.14\pm 7.38$ $132.33\pm 28.99$
Self-Consistency $8.65\pm 0.49$ $8.58\pm 0.76$ $34.20\pm 4.48$ $534.20\pm 124.48$
AP-Only $9.50\pm 0.35$ $8.18\pm 0.92$ $15.82\pm 3.13$ $122.50\pm 26.35$
TbS-Only $9.00\pm 0.00$ $8.43\pm 0.95$ $31.47\pm 12.61$ $141.50\pm 23.88$
CAST (AP+TbS) 9.67 $\pm$ 0.29 $8.52\pm 0.98$ $33.11\pm 10.18$ $122.51\pm 33.66$
D.S.-V3.2-Exp Zero-shot CoT $8.28\pm 0.87$ $8.78\pm 0.77$ $10.19\pm 4.47$ 29.36 $\pm$ 6.97
Few-shot CoT $8.36\pm 0.77$ $8.93\pm 0.70$ $7.64\pm 1.69$ 27.10 $\pm$ 12.53
Self-Consistency $8.51\pm 0.48$ $8.79\pm 0.82$ $24.24\pm 4.41$ $78.28\pm 33.17$
AP-Only $8.36\pm 0.97$ $8.90\pm 0.74$ 6.11 $\pm$ 0.65 $33.25\pm 12.17$
TbS-Only $9.50\pm 0.61$ $9.04\pm 0.67$ $11.51\pm 1.51$ $33.10\pm 10.62$
CAST (AP+TbS) $9.56\pm 0.26$ 9.14 $\pm$ 0.76 $14.07\pm 2.64$ $41.58\pm 17.73$
Claude-Sonnet-4.6 Zero-shot CoT $8.92\pm 0.56$ $9.21\pm 0.77$ $18.03\pm 3.74$ $10.30\pm 2.10$
CAST (AP+TbS) $9.07\pm 0.42$ $9.44\pm 0.60$ $20.26\pm 4.12$ $11.80\pm 2.50$
LLaMA-4-Maverick Zero-shot CoT $8.14\pm 0.87$ $8.21\pm 0.79$ $18.41\pm 4.47$ $25.40\pm 6.97$
CAST (AP+TbS) $8.29\pm 0.63$ $8.68\pm 0.76$ $21.43\pm 5.12$ $36.30\pm 8.73$